CaCO3>CO2 + CaO
How many grams of calcium carbonate will I need to form 3.45 liters of carbon dioxide?
1
Expert's answer
2013-02-18T10:29:34-0500
For answering this question you need to find the moles of CO2

& 3.45 n
CO2 -----> CO2
& 22.4 1

n=3.45*1 /& 22.4 = 0.154 mol
1 mol CO2 is produced from 1 mol CaCO3, so 0.154 mol of& CaCO3 can produce 0.154 mol of CO2.
The molar mass of CaCO3 is 100 g/mol.
m = n*Mw

m = 0.154 mol CaCO3 x (100 g / 1 mol ) =15.4 grams CaCO3